Hydrolysis of a compound B formula C7 H6cl2 was carried out by refluxing with excess potassium hydroxide solution.The resulting solution was acidified with dilute nitric acid and excess silver nitrate added to precipitate the chloride ions as silver chloride, 0.718 g of silver chloride was formed from 0.805 g of B.

a:Calculate the number of chlorine ions present in each molecule of B which are liberated by hydrolysis.

b:Suggest an explanation for the results obtained in a

c:Give a possible structure for B



1
Expert's answer
2021-05-11T03:22:01-0400

a) "mols AgCl =\frac {grams} {molar mass}"


"= \frac {0.718} {143.3} = 0.005"


"mols B = \frac {0.805} {161} = 0.005"


b) the AgNO3 shows just one molecule because only one atom Cl per molecule B was hydrolyzed.


c)